Shift report 20000816 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 16 2000 - 16:24:56 EDT

  • Next message: http account: "Shift report 20000816 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Aug 16 16:24:56 2000, Shift: 08:00-16:00
    Supervisor: sjs, with: sjs,hb
    Log book#4 pages 106 - 107
    
    Shift Summary:
    ------------------------------------------------------------
    1. Low intensity until noon. Good beam from 13:15-14:15,then
    beam lost and not recovered during shift.
    2. 2286 (started 7:20,stopped 10:10) MRS 90 deg, D5@0 kG
       2289 D5 @ -4kG (I=700A) (low beam intensity)
       2290 Camac pedestal, 2291 TPC pedestal
       2295 B/Y beams now 10418/6885, 24328 collisions
    3. No obvious problems...
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 16 2000 - 16:47:19 EDT